# WISE: A World Knowledge-Informed Semantic Evaluation for Text-to-Image Generation

### v1

@@ Abstract
Text-to-Image (T2I) models are capable of generating high-quality artistic creations and visual content. However, existing research and evaluation standards predominantly focus on image realism and shallow text-image alignment, lacking a comprehensive assessment of complex semantic understanding and world knowledge integration in text to image generation. To address this challenge, we propose WISE, the first benchmark specifically designed for World Knowledge-Informed Semantic Evaluation. WISE moves beyond simple word-pixel mapping by challenging models with 1000 meticulously crafted prompts across 25 sub-domains in cultural common sense, spatio-temporal reasoning, and natural science. To overcome the limitations of traditional CLIP metric, we introduce WiScore, a novel quantitative metric for assessing knowledge-image alignment. Through comprehensive testing of 20 models (10 dedicated T2I models and 10 unified multimodal models) using 1,000 structured prompts spanning 25 subdomains, our findings reveal significant limitations in their ability to effectively integrate and apply world knowledge during image generation, highlighting critical pathways for enhancing knowledge incorporation and application in next-generation T2I models. Code and data are available at PKU-YuanGroup/WISE.
@@ 1 Introduction
Text-to-Image (T2I) models [53] are capable of generating high-resolution and visually appealing images. However, these models often struggle with factual accuracy, particularly when presented with prompts requiring complex semantic understanding and world knowledge. This deficiency primarily arises from their limited incorporation of world knowledge [51] – the vast and diverse information, facts, and relationships that constitute real-world understanding. Critically, existing evaluation benchmarks fail to adequately assess this core capability, hindering progress towards truly robust and reliable T2I systems.
@@
Figure 1: Comparison of previous straightforward benchmarks and our proposed WISE. (a) Previous benchmarks typically use simple prompts, such as “A photo of two bananas” in GenEval [9], which only require shallow text-image alignment. (b) WISE, in contrast, uses prompts that demand world knowledge and reasoning, such as “Einstein’s favorite musical instrument,” to evaluate a model’s ability to generate images based on deeper understanding.
@@
Most T2I benchmarks suffer from a lack of semantic complexity. As shown in Figure 1, they use overly straightforward and simple prompts, failing to effectively challenge models’ ability to understand and generate images based on the model’s world knowledge. Furthermore, the most commonly used metric FID [12] primarily focuses on the realism of generated images. Although some benchmarks [11, 46, 50] utilize models like CLIP [30] to assess image-text semantic consistency scoring. However, CLIP’s limitations [52] in capturing fine-grained semantic information and handling complex reasoning hinder a comprehensive assessment of models’ performance in processing intricate semantic information. Consequently, existing evaluations fail to fully reveal the potential of models in real-world scenarios, particularly in tasks requiring world knowledge. For instance, when generating an image depicting a “tadpole that has undergone metamorphosis,” a model needs not only to comprehend the textual description (“tadpole,” “metamorphosis”) but also to invoke its internal world knowledge. This includes understanding amphibian development, the specific morphological changes involved (e.g., the growth of legs, the loss of the tail, the development of lungs), and the biological processes driving this transformation.
@@
To address the aforementioned problems, we propose a novel benchmark, WISE (World Knowledge-Informed Semantic Evaluation), designed to holistically assess models’ capabilities in semantic understanding and world knowledge integration through more complex T2I prompts. WISE covers three major domains: natural science, spatio-temporal reasoning, and cultural common sense, encompassing 25 sub-domains and a total of 1000 evaluation prompts. To rigorously assess the alignment of generated images with world knowledge, we introduce WiScore, a novel composite metric that emphasizes the accurate depiction of objects and entities within the generated image. WiScore is calculated as a weighted average of three key aspects: Consistency, Realism, and Aesthetic Quality.
@@
Moreover, traditional benchmarks always focus on dedicated T2I models, overlooking the potential of unified multimodal models, which integrate powerful LLMs trained on extensive text and image-text pairs and possess demonstrably strong world knowledge. While some studies have begun to explore whether the strong understanding capabilities of these unified models can benefit image generation, they often rely on overly simplistic benchmarks, thus failing to sufficiently prove this phenomenon. We broadened the scope of our evaluation beyond traditional dedicated T2I models. We employed our novel benchmark, WISE, to evaluate a total of 20 T2I models, encompassing both 10 dedicated T2I models and 10 unified multimodal models. However, experiment results demonstrate significant deficiencies in complex semantic understanding and world knowledge integration across existing T2I models. Even for unified multimodal models, their strong understanding capabilities do not fully translate into advantages in image generation, as revealed by our WISE evaluation. Despite their theoretical advantages, unified models generally underperform compared to dedicated T2I models in leveraging world knowledge for image generation. This indicates that current approaches to integrating LLMs within unified multimodal models may not yet fully unlock their potential for image generation that effectively integrates and applies world knowledge.
Our main contributions are as follows:
@@
Figure 3: Detailed composition of WISE, consisting of 3 categories and 25 subdomains.
@@ 2.1 Dedicated T2I Models
Dedicated T2I models represent the mainstream approach in the T2I field and have achieved remarkable progress in recent research. Currently, these models primarily fall into two categories: autoregressive models and diffusion models. Autoregressive [3, 6, 10, 40, 37, 28] models treat image generation as a sequence generation problem, similar to text generation. However, due to their computational cost and limitations in image quality, diffusion models have become the dominant paradigm. Diffusion [13, 32, 18, 34, 23] models iteratively add noise to an image and then progressively denoise it, often using pre-trained text encoders (e.g., CLIP [30]) to transform text prompts into embeddings that guide the denoising process. Key advancements include GLIDE [27], which pioneered diffusion models for T2I; Latent Diffusion Models (LDMs) [33], which improve quality and efficiency by operating in a latent space; and the Stable Diffusion series [29, 5], a landmark achievement built on LDMs.
@@ 2.2 Unified Multimodal Models
Unified multimodal models aim to construct general-purpose models capable of processing both textual and visual inputs, and performing cross-modal generation and understanding. These models [39, 16, 8, 43, 44, 25, 4, 48, 17, 45, 21, 38, 41, 35] are typically built upon powerful large language models (LLMs) [54] and extend next-token prediction [2] to image generation: the LLM generates visual tokens, and a VQ-VAE [42] or Diffusion model serves as a detokenizer. Moreover, Transfusion [55] and Show-O [49] demonstrate that bidirectional image diffusion can be combined with autoregressive text prediction within the same framework. D-DiT [22] achieves both Text-to-Image (T2I) and Image-to-Text (I2T) tasks using an end-to-end diffusion model. A crucial question concerning unified multimodal models is whether their understanding and generation capabilities can mutually enhance each other. Some studies [41, 45] have provided evidence supporting this phenomenon. However, in contrast to the rich and comprehensive benchmarks for multimodal understanding, T2I benchmarks are often relatively simple, lacking an in-depth examination of complex semantic understanding and world knowledge reasoning, making it difficult to fully prove the phenomenon.
@@ 2.3 Text to image evaluation
Despite the Fréchet Inception Distance (FID) [12] being one of the most widely adopted metrics for evaluating the quality of generated images, it falls short in assessing text-image consistency, thus failing to comprehensively measure the capabilities of text-to-image models. To address this deficiency, researchers have introduced a series of more sophisticated and challenging benchmarks [36, 24, 47] and evaluation metrics [11, 19, 24]. For instance, DPG-Bench [14] focuses on evaluating models’ ability in dense prompt following. T2I-CompBench [15] provides a benchmark suite for evaluating compositional generation, where prompts typically combine multiple distinct attributes. Furthermore, GenEval [9] serves as an object-centric evaluation framework specifically designed to assess compositional attributes of images, such as object co-occurrence, position, number, and color. However, the prompts used in these benchmarks are mostly straightforward, primarily examining models’ ability to follow compositional instructions for generation. Recently, some studies have shifted focus towards evaluating the application of specific types of knowledge in T2I models, such as physical reasoning in PhyBench [26] and broad common-sense knowledge in Commonsense-T2I [7]. Nevertheless, these emerging benchmarks are still very limited in their research scope and the quantity of evaluations.
3 Our benchmark: WISE
3.1 WISE construction
@@
Going beyond simply mapping words to pixels, we propose to evaluate world knowledge of current T2I models, which refers to the vast and diverse information, facts, and relationships that constitute our understanding of the real world. In our work, we focus on common world knowledge that can be represented visually. As shown in Figure 2, WISE comprises 1000 prompts designed to assess T2I models’ understanding and application of this knowledge across three major domains: Cultural Common Sense, Spatio-temporal Reasoning, and Natural Science, which would be divided into 25 subdomains.
@@
Data Collection and Prompt Design. We collected prompts from a variety of sources, including educational materials, encyclopedias, common knowledge question sets, and synthetic data generated by LLMs. These initial prompts were then refined and expanded by human annotators to ensure clarity, complexity, and unambiguous ground truth. Each prompt is accompanied by an explanation that clarifies the required world knowledge and the reasoning process needed for successful image generation. For example, a prompt like “The plant often gifted on Mother’s Day” would have an explanation like “The model should generate an image of a bouquet of carnations, a popular flower symbolizing love and admiration, often given on Mother’s Day.”
Table 1: Normalized WiScore of different models.

### v2

@@ Abstract
Text-to-Image (T2I) models are capable of generating high-quality artistic creations and visual content. However, existing research and evaluation standards predominantly focus on image realism and shallow text-image alignment, lacking a comprehensive assessment of complex semantic understanding and world knowledge integration in text-to-image generation. To address this challenge, we propose WISE, the first benchmark specifically designed for World Knowledge-Informed Semantic Evaluation. WISE moves beyond simple word-pixel mapping by challenging models with 1000 meticulously crafted prompts across 25 subdomains in cultural common sense, spatio-temporal reasoning, and natural science. To overcome the limitations of traditional CLIP metric, we introduce WiScore, a novel quantitative metric for assessing knowledge-image alignment. Through comprehensive testing of 22 models (10 dedicated T2I models and 12 unified multimodal models) using 1,000 structured prompts spanning 25 subdomains, our findings reveal significant limitations in their ability to effectively integrate and apply world knowledge during image generation, highlighting critical pathways for enhancing knowledge incorporation and application in next-generation T2I models. Code and data are available at PKU-YuanGroup/WISE.
@@ 1 Introduction
Text-to-image (T2I) models [53] are able to generate high-quality images that visually match explicit text descriptions. However, these models often have difficulty in ensuring true factual accuracy, especially when dealing with prompts that require complex semantic understanding (implicit understanding) and world knowledge (intrinsic knowledge matching). This shortcoming mainly stems from their limited ability to model the information behind world knowledge [51] - this information constitutes the massive facts and complex relationships required for real-world understanding. Although current unified multimodal models have begun to use LLMs’ powerful text modeling and implicit information extraction capabilities to break through this bottleneck through specific representations combined with generative model decoders [28; 41], existing evaluation benchmarks still have obvious limitations. They only focus on the surface alignment of images and texts, and fail to evaluate the core capabilities of T2I models from the perspective of implicit understanding and intrinsic knowledge matching, which seriously hinders the development of intelligent T2I systems.
@@
Figure 1: Comparison of previous straightforward benchmarks and our proposed WISE. (a) Previous benchmarks typically use simple prompts, such as "A photo of two bananas" in GenEval [9], which only require shallow text-image alignment. (b) WISE, in contrast, uses prompts that demand world knowledge and reasoning, such as "Einstein’s favorite musical instrument," to evaluate a model’s ability to generate images based on deeper understanding.
@@
Specifically, most T2I benchmarks suffer from a lack of semantic complexity. As shown in Figure 1, they use overly straightforward and simple prompts, failing to effectively challenge models’ ability to understand and generate images based on the model’s world knowledge. Furthermore, the most commonly used metric FID [12], primarily focuses on the realism of generated images. Although some benchmarks [11; 46; 50] utilize models like CLIP [31] to assess image-text semantic consistency scoring. However, CLIP’s limitations [52] in capturing fine-grained semantic information and handling complex reasoning hinder a comprehensive assessment of models’ performance in processing intricate semantic information. Consequently, existing evaluations fail to fully reveal the potential of models in real-world scenarios, particularly in tasks requiring world knowledge. For instance, when generating an image depicting a "tadpole that has undergone metamorphosis," a model needs not only to comprehend the textual description ("tadpole", "metamorphosis") but also to invoke its internal world knowledge. This includes understanding amphibian development, the specific morphological changes involved (e.g., the growth of legs, the loss of the tail, the development of lungs), and the biological processes driving this transformation.
@@
To address this problem, we propose a new benchmark WISE (World Knowledge-Informed Semantic Evaluation). This benchmark systematically evaluates the implicit semantic understanding and world knowledge integration capabilities of T2I models beyond basic text-image alignment through indirect textual cues. WISE covers three major areas: natural sciences, spatiotemporal reasoning, and cultural common sense, and contains 1,000 evaluation questions in 25 sub-areas. To rigorously evaluate the consistency of generated images with world knowledge, we design a novel comprehensive metric WiScore, which focuses on the accurate depiction of objects and entities under specific knowledge by weighted calculation of three key dimensions: consistency, realism, and aesthetic quality.
@@
Moreover, traditional benchmarks always focus on dedicated T2I models, overlooking the potential of unified multimodal models, which integrate powerful LLMs trained on extensive text and image-text pairs and possess demonstrably strong world knowledge. While some studies have begun to explore whether the strong understanding capabilities of these unified models can benefit image generation, they often rely on overly simplistic benchmarks, thus failing to sufficiently prove this phenomenon.
We broadened the scope of our evaluation beyond traditional dedicated T2I models. We employed our novel benchmark, WISE, to evaluate a total of 22 T2I models, encompassing both 10 dedicated T2I models and 12 unified multimodal models. However, experiment results demonstrate significant deficiencies in complex semantic understanding and world knowledge integration across existing T2I models. Even for unified multimodal models, their strong understanding capabilities do not fully translate into advantages in image generation, as revealed by our WISE evaluation. Despite their theoretical advantages, unified models generally underperform compared to dedicated T2I models in leveraging world knowledge for image generation. This indicates that current approaches to integrating LLMs within unified multimodal models may not yet fully unlock their potential for image generation that effectively integrates and applies world knowledge. Our main contributions are as follows:
@@ 2.1 Dedicated T2I Models
Dedicated T2I models represent the mainstream approach in the T2I field and have achieved remarkable progress in recent research. Currently, these models primarily fall into two categories: autoregressive models and diffusion models. Autoregressive [3; 6; 10; 40; 37; 29] models treat image generation as a sequence generation problem, similar to text generation. However, due to their computational cost and limitations in image quality, diffusion models have become the dominant paradigm. Diffusion [13; 32; 18; 34; 23] models iteratively add noise to an image and then progressively denoise it, often using pre-trained text encoders (e.g., CLIP [31]) to transform text prompts into embeddings that guide the denoising process. Key advancements include GLIDE [27], which pioneered diffusion models for T2I; Latent Diffusion Models (LDMs) [33], which improve quality and efficiency by operating in latent space; and Stable Diffusion series [30; 5], a landmark achievement built on LDMs.
@@ 2.2 Unified Multimodal Models
Unified multimodal models aim to construct general-purpose models capable of processing both textual and visual inputs, and performing cross-modal generation and understanding. These models [39; 16; 8; 43; 44; 25; 4; 48; 17; 45; 21; 38; 41; 35] are typically built upon powerful large language models (LLMs) [54] and extend next-token prediction [2] to image generation: the LLM generates visual tokens, and a VQ-VAE [42] or Diffusion model serves as a detokenizer. Moreover, Transfusion [55] and Show-O [49] demonstrate that bidirectional image diffusion can be combined with autoregressive text prediction within the same framework. D-DiT [22] achieves both Text-to-Image (T2I) and Image-to-Text (I2T) tasks using an end-to-end diffusion model. A crucial question concerning unified multimodal models is whether their understanding and generation capabilities can mutually enhance each other. Some studies [41; 45] have provided evidence supporting this phenomenon. However, in contrast to the rich and comprehensive benchmarks for multimodal understanding, T2I benchmarks are often relatively simple, lacking an in-depth examination of complex semantic understanding and world knowledge reasoning, making it difficult to fully prove the phenomenon.
@@ 2.3 Text to Image Evaluation
Despite the Fréchet Inception Distance (FID) [12] being one of the most widely adopted metrics for evaluating the quality of generated images, it falls short in assessing text-image consistency, thus failing to comprehensively measure the capabilities of text-to-image models. To address this deficiency, researchers have introduced a series of more sophisticated and challenging benchmarks [36; 24; 47] and evaluation metrics [11; 19; 24]. For instance, DPG-Bench [14] focuses on evaluating models’ ability in dense prompt following. T2I-CompBench [15] provides a benchmark suite for evaluating compositional generation, where prompts typically combine multiple distinct attributes. Furthermore, GenEval [9] serves as an object-centric evaluation framework specifically designed to assess compositional attributes of images, such as object co-occurrence, position, number, and color. However, the prompts used in these benchmarks are mostly straightforward, primarily examining models’ ability to follow compositional instructions for generation. Recently, some studies have shifted focus towards evaluating the application of specific types of knowledge in T2I models, such as physical reasoning in PhyBench [26] and broad common-sense knowledge in Commonsense-T2I [7]. Nevertheless, these emerging benchmarks are still very limited in their research scope and the quantity of evaluations.
3 The World Knowledge-Informed Semantic Evaluation (WISE) Benchmarks
Existing text-based image evaluation systems have two limitations: first, traditional benchmarks lack in-depth exploration of world knowledge; second, the current mainstream evaluation metrics for measuring image-text alignment only focus on the semantic level. Therefore, we will discuss in detail how to construct a world knowledge-informed T2I benchmark in subsection 3.1 and the necessity of establishing a robust metric by considering the bottlenecks of the current metric in subsection 3.2.
3.1 Building a Benchmark Based on World Knowledge
@@
Most existing benchmarks adopt prompt designs that are overly straightforward and lack semantic complexity. They primarily evaluate whether models can combine visual elements as explicitly instructed, but fail to assess the models’ capacity for complex semantic understanding and integration of world knowledge in text-to-image generation. Going beyond simply mapping words to pixels, we propose to evaluate world knowledge of current T2I models, which refers to the vast and diverse information, facts, and relationships that constitute our understanding of the real world. In our work, we focus on common world knowledge that can be represented visually. As shown in Figure 3, WISE comprises 1000 prompts designed to assess T2I models’ understanding and application of this knowledge across three major domains: Cultural Common Sense, Spatio-temporal Reasoning, and Natural Science, which are divided into 25 subdomains.
@@
Data Collection and Prompt Design. We collected prompts from a variety of sources, including educational materials, encyclopedias, common sense problem sets, and synthetic data generated by LLMs. These initial prompts were then refined and extended by human annotators to ensure their clarity, complexity, and explicit ground truth. Each prompt is accompanied by an explanation that describes the world knowledge and reasoning required for a potentially successful image generation. Specifically, the details of our three parts are as follows:
Cultural common sense. The Cultural Common Sense domain of WISE aims to evaluate the ability of models to understand specific cultural knowledge and apply it to image generation, which reflects the key aspects of understanding the real world. Models should not only understand the visual features of objects (e.g., shape, size, and color), but also match them with cultural knowledge of the real world. A model lacking such intrinsic knowledge matching will be an unintelligent image generator that can only grasp the surface correspondence between text nouns and visual elements, but cannot understand the cultural role of objects in the real world. This section covers a wide range of topics and is subdivided into 10 fine-grained sub-domains, including festivals, sports, religion, crafts, architecture, animals, plants, art, celebrities, and daily life. Together, these categories cover a wide range of cultural experiences and knowledge of humans. For example, prompts may involve generating images related to traditional festival customs, characteristic ethnic crafts, iconic landmarks, animals and plants with specific cultural significance, common sense in daily life, or events and objects related to celebrities.
Spatiotemporal reasoning. The spatiotemporal reasoning domain in WISE is structured around two key dimensions: temporal reasoning and spatial reasoning. Temporal reasoning is divided into Horizontal Temporal reasoning, which assesses understanding of relative temporal relationships between events or objects (e.g., "The Statue of Liberty at 10 pm Dubai time"), and Longitudinal Temporal reasoning, which assesses understanding of absolute temporal relationships, involving specific points in time (e.g., morning, noon, evening, season, specific year, or century). Spatial reasoning is divided into three subcategories: Different Views, which tests understanding of different perspectives, including top view, bottom view, side view, mirror image, and perspective effects; Geographic Relationships, which assesses understanding of spatial relationships between cities, countries, continents, and other geographic entities; and Relative Position, which focuses on understanding the spatial arrangement of objects in a scene relative to each other.
Natural sciences. Finally, the WISE benchmark includes a natural sciences domain that aims to assess whether models can not only understand scientific knowledge in a specific domain, but also use this understanding to reason about complex scientific scenarios and generate accurate and scientifically consistent images. At its core, generative models simulate the real world. Our goal is to assess whether these models’ understanding goes beyond mere visual replication and encompasses the complex science behind real-world phenomena. For example, a model should not only be able to generate images of water, ice, and steam, but also understand the underlying thermodynamic principles that govern transitions between these states (e.g., freezing, evaporation, condensation). This domain goes beyond general knowledge and delves into specialized bodies of knowledge in biology, physics, and chemistry.
3.2 Discovering the Deep Visual Language Alignment Bottlenecks of Traditional Evaluation
Existing metrics such as CLIP-Score [11] and VQA-Score [19] offer partial solutions for evaluating text-to-image alignment, but both fall short when faced with complex, knowledge-intensive prompts. CLIP-Score, constrained by CLIP’s bag-of-words limitation, struggles with compositional semantics and relational reasoning. VQA-Score improves on this by leveraging likelihood estimation via question answering, but still lacks sensitivity to implicit visual understanding and world knowledge grounding.To address these limitations, we propose a multi-faceted evaluation protocol to rigorously assess the quality of generated images, focusing on four key aspects: Consistency, Realism, Aesthetic Quality, and a composite metric, WiScore, shown in Figure 2.
Figure 4: Correlation between different metrics and human assessment.
Consistency evaluates the accuracy and completeness with which the generated image reflects the user’s prompt, capturing all key elements and nuances. Realism assesses the realism of the image, considering adherence to physical laws, accurate material representation, and coherent spatial relationships, determining how closely the image resembles a real photograph. Aesthetic Quality measures the overall artistic appeal and visual quality of the image, encompassing aspects such as composition, color harmony, and artistic style. WiScore, the central metric, emphasizes the accuracy of the depicted objects or entities within the generated image, directly reflecting our benchmark’s focus on world knowledge utilization. It is calculated as a weighted value of the other three metrics:
$\begin{gathered}\text{WiScore}=\frac{\alpha_{1}\times\text{Consistency}+\alpha_{2}\times\text{Realism}+\alpha_{3}\times\text{Aesthetic Quality}}{2},\\ \text{subject to }\alpha_{1}+\alpha_{2}+\alpha_{3}=1.\end{gathered}$ (1)
In this paper, we set $\alpha_{1}=0.7$, $\alpha_{2}=0.2$, and $\alpha_{3}=0.1$. This configuration not only prioritizes Consistency, reflecting the importance of accurately representing the prompt’s intended objects and their relationships, but also incorporates Realism and Aesthetic Quality to ensure overall image quality. Meanwhile, we observe that WiScore exhibits a strong positive correlation with human evaluations under certain hyperparameter settings, as detailed in Appendix B. A higher WiScore indicates superior performance in accurately depicting objects and concepts based on world knowledge. The detailed scoring criteria and complete results for each component score are provided in Appendix C. In the evaluation, we employ GPT-4o-2024-05-13, a powerful multimodal large language model, as the evaluator to assess the performance of T2I models. A carefully designed and rigorous scoring protocol is adopted to ensure consistency and reliability, with full implementation details provided in the Appendix D. As shown in Figure 4, among all automated metrics, WiScore exhibits the highest agreement with human judgments. The detailed comparison process can be found in the Appendix E.
4 Evaluation Results
4.1 Experiment Settings
We evaluate 22 T2I models, including 10 dedicated T2I models and 12 unified multimodal models. The dedicated T2I models are: stable-diffusion-v1-5 [33], stable-diffusion-2-1 [33], stable-diffusion-xl-base-0.9 [30], stable-diffusion-3-medium [5], stable-diffusion-3.5-medium [5], stable-diffusion-3.5-large [5], playground-v2.5-1024px-aesthetic [20], PixArt-XL-2-1024-MS [1], FLUX.1-dev [18], and FLUX.1-schnell [18]. The unified multimodal models are: GPT4o, Metaquery [28], Janus-Pro-7B [4], Janus-Pro-1B [4], JanusFlow-1.3B [25], Janus-1.3B [44], show-o [49], show-o-512 [49], Orthus-7B-instruct [17], vila-u-7b-256 [48], Liquid [45] and Emu3 [43]. For image generation, we use the official default configurations of each model and fix the random seed to ensure reproducibility. All experiments are conducted on eight NVIDIA A800 GPUs. For short, we use SD to denote stable-diffusion, playground-v2.5 for playground-v2.5-1024px-aesthetic, and PixArt-Alpha for PixArt-XL-2-1024-MS.
4.2 Main Results
Table 1: WiScore of different models. Based on six categories of basic knowledge, we calculated the results of different parameter models respectively, and the overall result was given as the weighted average of the previous parts to illustrate the ability of overall world knowledge. Bold indicates the best value in related items.